OpenStudy (anonymous):

The area of a circle is 28 in². The length of the radius is _____ in. 3.0 4.2 6.4 8.9

OpenStudy (anonymous):

\[Area = \pi r^{2}\] \[r = \sqrt{\frac{ Area }{ \pi }}\]

OpenStudy (anonymous):

28/3.14

OpenStudy (anonymous):

what did u get

OpenStudy (anonymous):

Yes and then the square root of that.

OpenStudy (anonymous):

3.0?

OpenStudy (anonymous):

yh 2.98 which is 3

OpenStudy (anonymous):

so i am correct